# arXiv:2603.02656v1[quant-ph] 03 Mar 2026

### 03 Mar 2026 v1

@@ Abstract
The graph isomorphism problem asks whether two graphs are identical up to vertex relabeling. While the exact problem admits quasi-polynomial-time classical algorithms, many applications in molecular comparison, noisy network analysis, and pattern recognition require a flexible notion of structural similarity. We study the quantum query complexity of approximate graph isomorphism testing, where two graphs on $n$ vertices drawn from the Erdős–Rényi distribution $\mathcal{G}(n,1/2)$ are considered approximately isomorphic if they can be made isomorphic by at most $k$ edge edits.
@@ 1.1 Motivation: Approximate Isomorphism
Many practical applications do not require exact isomorphism but rather a robust notion of structural similarity. In cheminformatics, comparing molecular graphs that may differ by a few bonds requires tolerance for small perturbations [RW02, WBD98]. In network analysis, noise in edge measurements makes exact isomorphism tests overly rigid [CFS+04]. In computer vision, pattern recognition under occlusion or measurement error demands flexible graph matching [BUN00, FSV01].
@@ 1.4 Related Work
Quantum walks have emerged as a powerful paradigm for quantum algorithm design. Discrete-time quantum walks on graphs were formalized by Aharonov, Ambainis, Kempe, and Vazirani [AAK+01], while Szegedy [SZE04] developed a general framework connecting classical Markov chains to quantum walks with quadratic speedup in hitting times. Magniez, Nayak, Roland, and Santha [MNR+11] extended this to quantum walk search, and Ambainis [AMB07] used quantum walks to achieve optimal query complexity for element distinctness.
@@
Classical algorithms for approximate or inexact graph matching include the work of Bunke and Shearer [BS98] on graph edit distance, Umeyama [UME88] on weighted graph matching via eigendecomposition, and recent work on sublinear-time algorithms for graph property testing [GOL17, FIS06]. The connection between graph similarity and spectrum has been extensively studied [vH03, WZ08].
@@
Quantum speedups for property testing problems have been demonstrated in various settings, including testing graph properties [ACR+10], testing Boolean functions [BFN+08], and distribution testing [BH11].
@@
Our work differs from the above lines of research in several respects. Unlike the hidden subgroup approach to exact GI, which requires quantum Fourier transforms over $\mathfrak{S}_{n}$ and faces representation-theoretic barriers [EHK04, MRŚ08], we work entirely in the adjacency-query model and target the easier approximate variant of the problem. Unlike classical sublinear graph property testing, where the natural random-sampling baseline is bottlenecked by the combinatorial explosion of the permutation search space, we exploit the MNRS quantum walk search framework [MNR+11] on a purpose-built compatibility product graph whose dense near-clique structure encodes approximate isomorphisms. This combination of quantum walk search on a structured product graph with Grover-accelerated reconstruction and verification is, to our knowledge, the first application of the MNRS framework to a graph similarity problem and yields a provable polynomial quantum speedup over any classical strategy.
@@ 2.2 Quantum Computation Basics
Theorem 2.7 (Amplitude estimation, [BHM+02]).
@@ 2.3 Quantum Walks
Theorem 2.10 (Quantum walk search, [MNR+11]).
@@ 3.1 Edit Distance and Approximate Isomorphism
Recall from Definition 2.4 that the edit distance $\operatorname{ed}(G,H)$ measures the minimum number of edge edits (additions and deletions) needed to make $G$ isomorphic to $H$, minimized over all vertex correspondences. We now define the decision problem.
@@ 6.2 Input Distributions
For any fixed permutation $\pi\in\mathfrak{S}_{n}$, the random variable $\operatorname{ed}_{\pi}(G,H)=\sum_{\{i,j\}}|(A_{G})_{ij}-(A_{H})_{\pi(i)\pi(j)}|$ is a sum of $\binom{n}{2}$ independent Bernoulli(1/2)(1/2) random variables (since $G$ and $H$ are independent Erdős–Rényi graphs). Hence $\mathbb{E}[\bar{\operatorname{ed}}_{\pi}(G,H)]=1/2$.
By Hoeffding’s inequality:
@@ 6.3 Indistinguishability Under Few Queries
For any deterministic adaptive algorithm making $q$ queries, the total variation distance between the distribution of query-answer transcripts under $\mathcal{D}_{\text{yes}}$ and $\mathcal{D}_{\text{no}}$ is at most
$d_{\mathrm{TV}}(\mathcal{T}_{\text{yes}},\mathcal{T}_{\text{no}})\;\leq\;\mathcal{O}\!\left(\frac{q^{2}}{n^{2}}\right).$
@@
We analyze the two distributions entry by entry. Under both distributions, $G$ is drawn from $\mathcal{G}(n,1/2)$, so queries to $A_{G}$ have identical distributions.
The key difference is in $H$:
  • Under $\mathcal{D}_{\text{yes}}$: $(A_{H})_{ij}=(A_{G})_{\pi^{-1}(i)\pi^{-1}(j)}\oplus B_{ij}$, where $\pi$ is a random permutation and $B_{ij}\sim\text{Bernoulli}(\varepsilon/2)$.
  • Under $\mathcal{D}_{\text{no}}$: $(A_{H})_{ij}$ is an independent Bernoulli(1/2)(1/2).
Consider the marginal distribution of the answer to a single query $(A_{H})_{ij}$. Under $\mathcal{D}_{\text{yes}}$, conditioned on $\pi$ and the queried entries of $A_{G}$:
$(A_{H})_{ij}=\begin{cases}(A_{G})_{\pi^{-1}(i)\pi^{-1}(j)}\oplus B_{ij}&\text{if }(\pi^{-1}(i),\pi^{-1}(j))\text{ has been queried in }G,\\ \text{Bernoulli}(1/2)&\text{otherwise}.\end{cases}$
[…]
The probability that a query $(A_{H})_{ij}$ “collides” with a previously queried entry in $G$ (via the hidden permutation $\pi$) is at most $q_{G}/(n^{2}-q_{G})$, where $q_{G}$ is the number of prior queries to $G$. Summing over all $q_{H}$ queries to $H$:
$\mathbb{E}[\text{number of collisions}]\;\leq\;\sum_{t=1}^{q_{H}}\frac{q_{G}}{n^{2}-t}\;\leq\;\frac{q_{G}\cdot q_{H}}{n^{2}-q}\;\leq\;\frac{q^{2}}{n^{2}}$
for $q\leq n^{2}/2$.
Each collision contributes at most $\varepsilon/2$ in total variation distance, and non-colliding queries have identical distributions. By the chain rule for total variation and the data processing inequality:
@@
$d_{\mathrm{TV}}(\mathcal{T}_{\text{yes}},\mathcal{T}_{\text{no}})\;\leq\;\frac{\varepsilon}{2}\cdot\frac{q^{2}}{n^{2}}\;\leq\;\mathcal{O}\!\left(\frac{q^{2}}{n^{2}}\right).$
@@ 6.4 Completing the Lower Bound Proof
Proof of Theorem 6.1.
By Yao’s minimax principle, it suffices to show that no deterministic algorithm making $q=o(n^{2})$ queries can distinguish $\mathcal{D}_{\text{yes}}$ from $\mathcal{D}_{\text{no}}$ with advantage >1/3>1/3.
By Lemmas 6.3 and 6.5, the distributions $\mathcal{D}_{\text{yes}}$ and $\mathcal{D}_{\text{no}}$ are supported on valid YES and NO instances, respectively, with probability $\geq 1-e^{-\Omega(n^{2})}$.
By Lemma 6.6, any deterministic algorithm making $q$ queries has distinguishing advantage at most $\mathcal{O}(q^{2}/n^{2})+e^{-\Omega(n^{2})}$.
For this to exceed $1/3$, we need $q^{2}/n^{2}=\Omega(1)$, i.e., $q=\Omega(n)$. However, the more careful analysis using the hidden permutation structure gives a tighter bound.
Claim 6.7.
The total variation bound can be tightened to $d_{\mathrm{TV}}(\mathcal{T}_{\text{yes}},\mathcal{T}_{\text{no}})\leq\mathcal{O}(q^{2}/n^{4}+q/n^{2})$.
Proof of Claim 6.7.
Warm-up: $\Omega(n)$ bound.  The bound $d_{\mathrm{TV}}\leq\mathcal{O}(q^{2}/n^{2})$ from Lemma 6.6 immediately gives $q=\Omega(n)$. We now strengthen this in two steps.
Warm-up: $\Omega(n^{4/3})$ via information theory.  The mutual information between the label $Y\in\{\text{yes},\text{no}\}$ and the transcript $\mathcal{T}$ is bounded by the chain rule:
$I(Y;\mathcal{T})\;\leq\;\sum_{t=1}^{q}I(Y;A_{t}\mid A_{1},\ldots,A_{t-1})\;\leq\;\sum_{t=1}^{q}\mathcal{O}\!\left(\frac{t}{n^{2}}\right)^{2}=\mathcal{O}\!\left(\frac{q^{3}}{n^{4}}\right).$
By Fano’s inequality, reliable distinguishing (error $\leq 1/3$) requires $I(Y;\mathcal{T})\geq\Omega(1)$, giving $q=\Omega(n^{4/3})$.
Tight bound: $\Omega(n^{2})$ via conditional independence.  We prove the $\Omega(n^{2})$ bound by a formal induction on the transcript.
Define the filtration $\mathcal{F}_{t}$ as the $\sigma$-algebra generated by the first $t$ query-answer pairs $(q_{1},a_{1}),\ldots,(q_{t},a_{t})$, where each $q_{\ell}$ specifies a matrix ($G$ or $H$) and an index pair, and $a_{\ell}\in\{0,1\}$ is the answer.
Subclaim 6.7.1.
@@
For every $t\in[q]$, conditioned on $\mathcal{F}_{t-1}$ and on the event that step $t$ does not produce a collision (i.e., the queried $H$-entry $(A_{H})_{ij}$ does not correspond, via $\pi$, to any previously queried $G$-entry), the conditional distribution of the answer $a_{t}$ is $\mathrm{Bernoulli}(1/2)$ under both $\mathcal{D}_{\text{yes}}$ and $\mathcal{D}_{\text{no}}$.
Proof of Subclaim 6.7.1.
@@
We argue by induction on $t$. The base case $t=1$ is immediate since any single entry of $A_{G}$ or $A_{H}$ is marginally $\mathrm{Bernoulli}(1/2)$ under both distributions.
For the inductive step, suppose the claim holds for all steps <t<t. At step $t$, the algorithm queries some entry. There are two cases:

### 28 Mar 2026 v2

# Quantum Algorithms for Approximate Graph Isomorphism Testing

@@ Abstract
The graph isomorphism problem asks whether two graphs are identical up to vertex relabeling. While the exact problem admits quasi-polynomial-time classical algorithms, many applications in molecular comparison, noisy network analysis, and pattern recognition require a flexible notion of structural similarity. We study the quantum query complexity of approximate graph isomorphism testing, where the two input graphs are marginally distributed as $\mathcal{G}(n,1/2)$ but may be arbitrarily correlated through a latent planted permutation, and are considered approximately isomorphic if they can be made isomorphic by at most $k$ edge edits.
@@ 1.1 Motivation: Approximate Isomorphism
Many practical applications do not require exact isomorphism but rather a robust notion of structural similarity. In cheminformatics, comparing molecular graphs that may differ by a few bonds requires tolerance for small perturbations [RW02, WBD98]. In network analysis, noise in edge measurements makes exact isomorphism tests overly rigid [CFSV04]. In computer vision, pattern recognition under occlusion or measurement error demands flexible graph matching [Bun00, FSV01].
The correlated-input model and why it is the right formulation.
A critical subtlety concerns the joint distribution of the input pair $(G,H)$. Two independent draws from $\mathcal{G}(n,1/2)$ satisfy $\bar{ed}(G,H)\geq 1/4$ with probability $1-e^{-\Omega(n^{2})}$ (Lemma 6.5), so an algorithm that outputs No without making a single query already succeeds with overwhelming probability. The NO case is therefore trivially easy under the independent model, and any nontrivial separation must come from something else. The meaningful and natural formulation—the one we adopt throughout—considers pairs $(G,H)$ whose marginal distributions are both $\mathcal{G}(n,1/2)$, yet which may be arbitrarily correlated:
  • YES instances: $G\sim\mathcal{G}(n,1/2)$; a permutation $\pi\sim\mathrm{Uniform}(S_{n})$ is drawn independently; then $H$ is obtained from $\pi(G)$ by independently flipping each edge with probability $\varepsilon/2$. Marginally, $H\sim\mathcal{G}(n,1/2)$, but $(G,H)$ are strongly correlated through the latent $\pi$.
  • NO instances: $G$ and $H$ are drawn independently from $\mathcal{G}(n,1/2)$, so $(G,H)$ are uncorrelated and $\bar{ed}(G,H)\geq 2\varepsilon$ w.o.p.
From the marginals alone the two cases are identical; an algorithm must probe the joint adjacency structure to detect the latent correlation. This “correlated planted model” is standard in the graph alignment and hypothesis-testing literatures [PG11, CK16, MWZ24, RS21], and captures practical scenarios such as comparing two social-network snapshots of the same population observed at different times, or two molecular graphs that are noisy measurements of the same underlying molecule. Our upper bound (Theorem 5.9) and classical lower bound (Theorem 6.1) are both about the complexity of detecting this correlation, making the formulation internally consistent: the classical algorithm must work against a YES distribution that is genuinely hard, not one that is trivially easy from the NO side.
@@ 1.4 Related Work
Quantum walks have emerged as a powerful paradigm for quantum algorithm design. Discrete-time quantum walks on graphs were formalized by Aharonov, Ambainis, Kempe, and Vazirani [AAKV01], while Szegedy [Sze04] developed a general framework connecting classical Markov chains to quantum walks with quadratic speedup in hitting times. Magniez, Nayak, Roland, and Santha [MNRS11] extended this to quantum walk search, and Ambainis [Amb07] used quantum walks to achieve optimal query complexity for element distinctness.
@@
Classical algorithms for approximate or inexact graph matching include the work of Bunke and Shearer [BS98] on graph edit distance, Umeyama [Ume88] on weighted graph matching via eigendecomposition, and recent work on sublinear-time algorithms for graph property testing [Gol17, Fis06]. The connection between graph similarity and spectrum has been extensively studied [vDH03, WZ08].
@@
Quantum speedups for property testing problems have been demonstrated in various settings, including testing graph properties [ACR+10], testing Boolean functions [BFNR08], and distribution testing [BH11].
@@
Our work differs from the above lines of research in several respects. Unlike the hidden subgroup approach to exact GI, which requires quantum Fourier transforms over $\mathfrak{S}_{n}$ and faces representation-theoretic barriers [EHK04, MRŚ08], we work entirely in the adjacency-query model and target the easier approximate variant of the problem. Unlike classical sublinear graph property testing, where the natural random-sampling baseline is bottlenecked by the combinatorial explosion of the permutation search space, we exploit the MNRS quantum walk search framework [MNRS11] on a purpose-built compatibility product graph whose dense near-clique structure encodes approximate isomorphisms. This combination of quantum walk search on a structured product graph with Grover-accelerated reconstruction and verification is, to our knowledge, the first application of the MNRS framework to a graph similarity problem and yields a provable polynomial quantum speedup over any classical strategy.
@@ 2.2 Quantum Computation Basics
Theorem 2.7 (Amplitude estimation, [BHMT02]).
@@ 2.3 Quantum Walks
Theorem 2.10 (Quantum walk search, [MNRS11]).
@@ 3.1 Edit Distance and Approximate Isomorphism
Recall from Definition 2.4 that $\mathrm{ed}(G,H)$ measures the minimum number of edge edits needed to make $G$ isomorphic to $H$. As argued in Section 1.1, the meaningful decision problem is one in which both graphs are marginally $\mathcal{G}(n,1/2)$ but may be arbitrarily correlated; the algorithm must detect the presence or absence of a latent planted permutation.
@@
Remark 3.2 (Explicit correlation structure of the input distributions).
To make the correlation structure completely explicit: in the YES case the joint distribution of $(G,H)$ is
$G\sim\mathcal{G}(n,\tfrac{1}{2}),\pi\sim\mathrm{Uniform}(S_{n})\text{ independent of }G,(A_{H})_{ij}=(A_{G})_{\pi^{-1}(i)\pi^{-1}(j)}\oplus B_{ij},B_{ij}\overset{\mathrm{iid}}{\sim}\mathrm{Bern}(\varepsilon/2),$
so that marginally $H\sim\mathcal{G}(n,\tfrac{1}{2})$. In the NO case $G$ and $H$ are drawn independently from $\mathcal{G}(n,\tfrac{1}{2})$, giving the same marginals but zero correlation. An algorithm that has access only to the marginals of $G$ and $H$ individually cannot distinguish the two cases; it must compare the two adjacency matrices jointly.
@@ 6.2 Input Distributions
For any fixed permutation $\pi\in\mathfrak{S}_{n}$, the random variable $\operatorname{ed}_{\pi}(G,H)=\sum_{\{i,j\}}|(A_{G})_{ij}-(A_{H})_{\pi(i)\pi(j)}|$ is a sum of $\binom{n}{2}$ independent Bernoulli(1/2)(1/2) random variables (since $G$ and $H$ are independent Erdős–Rényi graphs). Hence $\mathbb{E}[\bar{\operatorname{ed}}_{\pi}(G,H)]=1/2$. By Hoeffding’s inequality:
@@ 6.3 Indistinguishability Under Few Queries
Let $\mathcal{A}$ be any deterministic adaptive algorithm making $q$ queries in total, of which $q_{G}$ go to $A_{G}$ and $q_{H}=q-q_{G}$ go to $A_{H}$. Let $T_{\mathrm{yes}}$ and $T_{\mathrm{no}}$ denote the distributions over query-answer transcripts under $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$ respectively. Provided $q\leq\binom{n}{2}/2$,
$d_{\mathrm{TV}}\!\left(T_{\mathrm{yes}},\,T_{\mathrm{no}}\right)\;\leq\;\frac{\varepsilon\,q_{G}\,q_{H}}{\binom{n}{2}-q}\;\leq\;\frac{4\varepsilon q^{2}}{n^{2}}.$
@@
Write the transcript as a sequence of random variables $\mathcal{T}=(Q_{1},A_{1},\ldots,Q_{t},A_{t},\ldots,Q_{q},A_{q})$, where $Q_{t}$ specifies which matrix ($G$ or $H$) and which entry to query, and $A_{t}\in\{0,1\}$ is the answer. Because $\mathcal{A}$ is deterministic and adaptive, $Q_{t}$ is a deterministic function of $(A_{1},\ldots,A_{t-1})$.
Chain rule for total variation.
By the chain rule for total variation (applied iteratively to conditional distributions),
$d_{\mathrm{TV}}(T_{\mathrm{yes}},T_{\mathrm{no}})\;\leq\;\sum_{t=1}^{q}\mathbb{E}\!\left[d_{\mathrm{TV}}\!\left(\mathcal{L}(A_{t}\mid Q_{t},\mathcal{F}_{t-1})_{\mathrm{yes}},\;\mathcal{L}(A_{t}\mid Q_{t},\mathcal{F}_{t-1})_{\mathrm{no}}\right)\right],$ (2)
where $\mathcal{F}_{t-1}=(Q_{1},A_{1},\ldots,Q_{t-1},A_{t-1})$ is the $\sigma$-algebra of the first t1t-1 steps, and the expectation is over the joint randomness of the input and the algorithm up to step t1t-1. (This follows from the tensorization inequality $d_{\mathrm{TV}}(P_{1:q},Q_{1:q})\leq\sum_{t}\mathbb{E}_{P_{1:t-1}}[d_{\mathrm{TV}}(P_{t}(\cdot\mid\mathcal{F}_{t-1}),Q_{t}(\cdot\mid\mathcal{F}_{t-1}))]$, a standard consequence of the data-processing inequality applied to each step; see, e.g., [Duc19].)
Queries to $A_{G}$.
Under both $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$, $G\sim\mathcal{G}(n,\frac{1}{2})$ independently of everything else, so a previously unqueried entry $(A_{G})_{ij}$ is $\mathrm{Bernoulli}(\frac{1}{2})$, independent of $\mathcal{F}_{t-1}$, under both distributions. The per-step contribution to (2) is therefore 0 for all queries to $A_{G}$.
Queries to $A_{H}$: no collision.
Suppose step $t$ queries $(A_{H})_{ij}$. Define a collision at step $t$ as the event that the entry $(A_{G})_{\pi^{-1}(i)\pi^{-1}(j)}$ appears among the (at most $q_{G}^{(t)}\leq q_{G}$) positions of $A_{G}$ already queried. If no collision occurs at step $t$, then under $D_{\mathrm{yes}}$:
$(A_{H})_{ij}\;=\;\underbrace{(A_{G})_{\pi^{-1}(i)\pi^{-1}(j)}}_{\text{not yet queried}}\oplus B_{ij},$
where $B_{ij}\sim\mathrm{Bern}(\varepsilon/2)$ is independent of $\mathcal{F}_{t-1}$ and of $G$. Since $\pi$ is uniformly random and independent of $G$, the position $(\pi^{-1}(i),\pi^{-1}(j))$ is uniformly distributed over $\binom{[n]}{2}$; conditioned on no collision, it is uniform over the at-least $\binom{n}{2}-q_{G}$ unqueried positions of $A_{G}$. Any such entry is an independent $\mathrm{Bern}(\frac{1}{2})$ variable (by the Erdős–Rényi structure), also independent of $\mathcal{F}_{t-1}$. XOR-ing with the independent $B_{ij}\sim\mathrm{Bern}(\varepsilon/2)$ gives
$(A_{H})_{ij}\;\sim\;\mathrm{Bern}\!\left(\tfrac{1}{2}(1-\tfrac{\varepsilon}{2})+\tfrac{1}{2}\cdot\tfrac{\varepsilon}{2}\right)\;=\;\mathrm{Bern}(\tfrac{1}{2}).$
Under $D_{\mathrm{no}}$, $(A_{H})_{ij}\sim\mathrm{Bern}(\frac{1}{2})$ independently of $\mathcal{F}_{t-1}$ by construction. Hence the per-step contribution is 0 conditioned on no collision.
Queries to $A_{H}$: collision.
If a collision does occur at step tt—meaning $(A_{G})_{\pi^{-1}(i)\pi^{-1}(j)}$ was queried at some earlier step $s<t$ and revealed value $a_{s}\in\{0,1\}$—then under $D_{\mathrm{yes}}$:
@@
$(A_{H})_{ij}=a_{s}\oplus B_{ij}\;\sim\;\mathrm{Bern}\!\left(\tfrac{\varepsilon}{2}\right)\text{ if }a_{s}=0,\quad\mathrm{Bern}\!\left(1-\tfrac{\varepsilon}{2}\right)\text{ if }a_{s}=1.$
Under $D_{\mathrm{no}}$, $(A_{H})_{ij}\sim\mathrm{Bern}(\frac{1}{2})$. In both sub-cases,
$d_{\mathrm{TV}}\!\left(\mathrm{Bern}\!\left(\tfrac{1}{2}\pm\tfrac{\varepsilon}{2}\right),\,\mathrm{Bern}\!\left(\tfrac{1}{2}\right)\right)\;=\;\tfrac{\varepsilon}{2}.$
Bounding the collision probability per step.
At step $t$, there are at most $q_{G}^{(t)}\leq q_{G}$ previously queried positions of $A_{G}$. Conditioned on $\mathcal{F}_{t-1}$, the random position $(\pi^{-1}(i),\pi^{-1}(j))$ is uniform over $\binom{[n]}{2}$, so
$\Pr\!\left[\text{collision at step }t\;\middle|\;\mathcal{F}_{t-1}\right]\;\leq\;\frac{q_{G}^{(t)}}{\binom{n}{2}}\;\leq\;\frac{q_{G}}{\binom{n}{2}}$.
Assembling the bound.
Substituting into (2) and summing over the $q_{H}$ queries to $H$:
$d_{\mathrm{TV}}(T_{\mathrm{yes}},T_{\mathrm{no}})\;\leq\;\sum_{t:\,\text{query to }H}\Pr[\text{collision at }t]\cdot\frac{\varepsilon}{2}\;\leq\;q_{H}\cdot\frac{q_{G}}{\binom{n}{2}}\cdot\frac{\varepsilon}{2}\;=\;\frac{\varepsilon\,q_{G}\,q_{H}}{\binom{n}{2}}$.
Since $q_{G},q_{H}\leq q$ and $\binom{n}{2}\geq n^{2}/4$ for $n\geq 2$, and $\binom{n}{2}-q\geq\binom{n}{2}/2\geq n^{2}/8$ for $q\leq\binom{n}{2}/2$, we obtain
$d_{\mathrm{TV}}(T_{\mathrm{yes}},T_{\mathrm{no}})\;\leq\;\frac{\varepsilon q^{2}}{\binom{n}{2}}\;\leq\;\frac{4\varepsilon q^{2}}{n^{2}}.\qed$
@@ 6.4 Completing the Lower Bound Proof
We prove Theorem 6.1 in two parts. Part 1 establishes an $\Omega(n)$ lower bound via a total-variation argument on the random distributions $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$. Part 2 lifts this to $\Omega(n^{2})$ via an adaptive-query covering argument that applies to worst-case inputs.
Subclaim 6.6.1 (Independence on no collision).
@@
For every $t\in[q]$, conditioned on $\mathcal{F}_{t-1}$ and on the event $E_{t}$ that step $t$ does not produce a collision, the conditional distribution of the answer $A_{t}$ is $\mathrm{Bern}(\tfrac{1}{2})$ under both $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$.
Proof.
By induction on $t$.
@@
Base case ($t=1$). Any single entry of $A_{G}$ or $A_{H}$ is marginally $\mathrm{Bern}(\tfrac{1}{2})$ under both distributions, and there is no prior history, so the claim holds trivially.
Inductive step. Assume the claim holds for all steps <t<t.